# Understanding oscillating features of the time-like nucleon electromagnetic form factors within the extending vector meson dominance model

### 08 Dec 2023 v1

@@ I Introduction
Nucleons (proton and neutron) are composite particles consisting of three valence quarks ($uud$ and $udd$) and a neutral sea of strong interaction. Despite the discovery of the proton over 100 years ago, an exact theoretical description of its internal structure has not been achieved within the framework of quantum chromodynamics (QCD) theory. This is due to the non-perturbative nature of QCD in the energy regime of the nucleon. On the theoretical side, the nucleon electromagnetic structure can be described by the electromagnetic form factors (EMFFs), which depends on squared of the four-momentum ($q^{2}$) of the exchanged virtual photon. The EMFFs in the space-like region ($q^{2}<0$) are real and they can be associated with the charge and magnetic distribution of the nucleons Pacetti et al. (2015). However, in the time-like region ($q^{2}>0$), the EMFFs become complex, providing information on the time evolution of the nucleon’s charge and magnetic moments at the nucleon-antinucleon pair formation point  Belushkin et al. (2007); Kuraev et al. (2012); Bianconi and Tomasi-Gustafsson (2015); Tomasi-Gustafsson et al. (2021). On the experimental side, the EMFFs of nucleons were measured by both the $e^{-}N\to e^{-}N$ elastic scattering Gayou et al. (2002, 2001); Andivahis et al. (1994); Bosted et al. (1992); Jones et al. (2000) and $e^{+}e^{-}\to N\bar{N}$ annihilation reactions Lees et al. (2013); Achasov et al. (2014); Ablikim et al. (2015); Akhmetshin et al. (2016); Ablikim et al. (2019a, 2020a, 2021a, 2021b); Huang and Ferroli (2021); Achasov et al. (2022). One can obtain the space-like form factors from the former process and the time-like form factors from the latter process, respectively.
@@
A comprehensive analysis of the nucleon’s electromagnetic form factors in both the space- and time-like regions using dispersion theory is conducted in Ref. Lin et al. (2022). For the periodic behaviour of the nucleon’s electromagnetic form factors, it has been studied in Refs. Cao et al. (2022); Xia et al. (2022); Dai et al. (2022); Bianconi and Tomasi-Gustafsson (2022); Yang et al. (2022); Milstein and Salnikov (2022); Tomasi-Gustafsson and Pacetti (2022); Qian et al. (2023); Dai et al. (2023), using various phenomenological methods. It is worthy mentioning that the periodic behaviour of the nucleon’s electromagnetic form factors are induced by those broad vector mesons according to the studies in Ref. Cao et al. (2022). However, these vector mesons couple to $N\bar{N}$ pair very weekly. In fact, a unified description of the time- and space-like electromagnetic form factors of nucleons was proposed in Ref. Lomon and Pacetti (2012), where these low-lying vector mesons were taken into account. Meanwhile, in the time-like region, a revised Breit-Wigner formulas with momentum-dependent widths are needed for these vector resonances Lomon and Pacetti (2012). Indeed, the contributions of these vector mesons are important to the reactions of $e^{+}e^{-}$ annihilation into light hadrons Ablikim et al. (2019b, 2021c, 2021d, 2021e, c, 2020b, d, 2020c, 2021f).
@@
with $\gamma=1.41$ $\rm GeV^{-2}$ for the case of nucleon.
On the other hand, the BESIII Collaboration has also provided a large number of experimental data about $e^{+}e^{-}$ annihilation into light mesons Ablikim et al. (2019b, 2021c, 2021d, 2021e, c, 2020b, d, 2020c, 2021f), which provides the basis for our construction of light vector states. From the analyses of the $e^{+}e^{-}\to\pi^{+}\pi^{-}$, $\omega\pi^{0}$, and $\rho^{0}\eta^{\prime}$ reactions, these $\rho$ excited sates were studied in Refs. Wang et al. (2020); Zhou et al. (2022); Wang et al. (2021). The spectrum of excited $\rho$, $\omega$, and $\phi$ states were also investigated in Ref. Wang et al. (2022) using the modified Godfrey-Isgur model Song et al. (2015). By performing the phenomenological analysis, the two-body strong decay of the excited $\rho$ states are systematically studied He et al. (2013). Based on the above theoretical calculations and the vector mesons listed in the particle data group (PDG) Workman et al. (2022), we take $\rho(2D)$, $\omega(3D)$, and $\omega(5S)$ into account in this work. Their masses and widths are collected in Table 1. By considering the contributions of these above low-lying vector meson excited states, we studied the $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions within the VMD model. It is found that the total cross sections of the two above reactions (corresponding to the nucleon’s effective form factors) and the nucleon electromagnetic form factors can be well described. However, the so-called oscillation behavior of the nucleon’s EMFFs discovered by the BESIII Collaboration Ablikim et al. (2021b) is not confirmed.
@@
This article is organized as follows: in the next section, we show the theoretical formalism of the nucleon’s EMFFs in the VMD model. Numerical results of the nucleon’s effective form factors, the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions, and the oscillating features of the nucleon’s EMFFs are shown in Sec. III. A short summary is given in the last section.
@@ II Theoretical formalism
with $\alpha=e^{2}/(4\pi)$ the fine-structure constant and $\beta=\sqrt{1-4m_{N}^{2}/s}$ the phase-space factor, $m_{N}$ is the nucleon mass. Here, $C_{N}$ represents the S-wave Sommerfeld-Gamow factor accounts for the electromagnetic interaction of charged pointlike fermion pairs in the final state Arbuzov and Kopylova (2012). For proton, $C_{p}=y/(1-e^{-y})$ with $y=\frac{\alpha\pi}{\beta}\frac{2M_{p}}{\sqrt{s}}$, while for the neutron, $C_{n}=1$. Considering the $C_{N}$ factor, it is expected that the total cross section of $e^{+}e^{-}\to p\bar{p}$ reaction is nonzero at the reaction threshold.
@@
where $\tau=q^{2}/(4m_{N}^{2})$ and $q^{2}=s$ the invariant mass square of the $e^{+}e^{-}$ annihilation process. Once we have $F_{1}(s)$ and $F_{2}(s)$, we can naturally obtain the electromagnetic form factors of the nucleon. Then the total cross sections of $e^{+}e^{-}\to N\bar{N}$ can be easily calculated.
Because proton and neutron are isospin doublets, the $p\bar{p}$ and $n\bar{n}$ states are expressed in terms of isospin 0 and 1 components. The mixtures of isovector and isoscalar for $p\bar{p}$ and $n\bar{n}$ of equal relative weight but different signs are imposed by the isospin symmetry as introduced by the underlying Clebsch-Gordan coefficients. As discussed above, in addition to these ground isovector $\rho$ meson and isoscalar $\omega$ meson, we also consider the excited vector mesons with masses around 2.0 to 3.0 GeV.  [^1] It’s worth noting that due to the Okuba-Zweig-Iizuka rule, we do not consider the $\phi$ meson and its excited states here.
Then, the Dirac and Pauli form factors $F_{1}$ and $F_{2}$ for nucleon in the time-like region can be parameterized as,
@@
$B_{\rho}$ $=$ $\frac{m_{\rho}^{2}}{m_{\rho}^{2}-s-im_{\rho}\Gamma_{\rho}},$ (10)
$B_{\omega}$ $=$ $\frac{m_{\omega}^{2}}{m_{\omega}^{2}-s-im_{\omega}\Gamma_{\omega}},$ (11)
$B_{\rho(2D)}$ $=$ $\frac{m_{\rho(2D)}^{2}}{m_{\rho(2D)}^{2}-s-im_{\rho(2D)}\Gamma_{\rho(2D)}},$ (12)
$B_{\omega(3D)}$ $=$ $\frac{m_{\omega(3D)}^{2}}{m_{\omega(3D)}^{2}-s-im_{\omega(3D)}\Gamma_{\omega(3D)}},$ (13)
$B_{\omega(5S)}$ $=$ $\frac{m_{\omega(5S)}^{2}}{m_{\omega(5S)}^{2}-s-im_{\omega(5S)}\Gamma_{\omega(5S)}}.$ (14)
Here, we take $m_{\rho}=775.26$ MeV, $\Gamma_{\rho}=147.4$ MeV, $m_{\omega}=782.66$ MeV, and $\Gamma_{\omega}=8.68$ MeV. In addition, at $q^{2}=0$, with the constraints $G_{E}^{n}=0$ and $G_{M}^{n}=\mu_{n}$, $G_{E}^{p}=1$ and $G_{M}^{p}=\mu_{p}$, the coefficients $f_{1}^{n}$, $f_{2}^{n}$, $f_{1}^{p}$ and $f_{2}^{p}$ can be calculated,
@@ III NUMERICAL RESULTS
The results of the fitted parameters are listed in Table 2. The corresponding $\chi^{2}/\rm d.o.f$ is 1.6, where ${\rm d.o.f}$ is the number of dimension of the freedom.
@@
We show firstly the fitted numerical results of the electromagnetic form factors of proton and neutron in Fig. LABEL:fig:EMFFs, comparing with the experimental data of the BESIII Collaboration. Within the contributions of the excited $\rho(2D)$, $\omega(3D)$, and $\omega(5S)$ states, we can obtain a good description of the electromagnetic form factors for both proton and neutron. For the electric form factor of proton, there is a dip structure around $2.3$ GeV which is because of the interference between the contributions of $\omega(3D)$ and $\omega(5S)$, and there is a clear peak for the $\omega(5S)$. It is expected that these theoretical results can be tested by future experiments.
@@
Second, we show the fitted results of the modulus of the ratio $G_{E}/G_{M}$ for the proton in Fig. 2. The experimental data have large errors. Again, there are dip and bump structures because of the contributions of of $\omega(3D)$ and $\omega(5S)$ states. It is worthy mentioning that the ratio of $G_{E}/G_{M}$ equals to one at the reaction threshold as it should be. Clearly, more precise experimental data are needed to check our model calculations.
Figure 2: The ratio $|G_{E}/G_{M}|$ of proton.
@@
The fitted numerical results of the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions are shown in Figs. 3 (a) and (d), respectively. One can see that, thanks to the contribution of $\rho(2D)$, the platform behavior of the total cross section of $e^{+}e^{-}\to p\bar{p}$ can be well reproduced, and the experimental data of the $e^{+}e^{-}\to n\bar{n}$ can also be described within the fitted model parameters. Furthermore, these nonmonotonic structures around $2.2$ to $2.4$ GeV can be fairly well reproduced by considering the contributions of $\omega(3D)$ and $\omega(5S)$ states. In fact, a resonance with mass around 2300 MeV and width about 188 MeV is needed in the analysis of the total cross sections of $e^{+}e^{-}\to p\bar{p}$ reaction in Ref. Xia et al. (2022), where a simple Breit-Wigner amplitude for the $e^{+}e^{-}\to p\bar{p}$ reaction was taken.
@@
Since we can reproduce the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions, it is expected that the effective form factors of proton and neutron can also be described well, as shown in Figs. 3 (b) and (e), respectively. This indicates that the nonmonotonic line shapes of the nucleon effective form factors can be explained within the VMD model, where the contributions of the excited vector mesons are taken into account.
@@
Next, we will discuss the nonmonotonic features of the nucleon effective form factors. In general, the main part of the effective form factors of the nucleon can be descried by a dipole form as shown in Eq. (1) with a global factor $c_{0}$, $G_{D}=c_{0}/(1-\gamma s)^{2}$. By fitting the experimental data of the effective form factors of proton and neutron, we get $c_{0}=5.54\pm 0.02$ for proton and $c_{0}=4.08\pm 0.04$ for neutron. Then, subtracting the obtained effective form factor $|G_{\rm eff}|$ from $G_{\rm D}$, one can get the residual $G_{\rm osc}$,
@@
The predictions of our model for the oscillation parts of the nucleon effective form factors are shown in Figs. 3 (c) and (f), respectively. It is found that the model predictions are in agreement with the subtracted data obtained with Eq. (19).
@@
Note that, in Ref. Ablikim et al. (2021b), to get the conclusion which the effective form factor of both proton and neutron show a periodic behaviour, different formula of $G_{D}$ were taken for proton and neutron. A three-pole formula $F_{3p}=F_{0}/[(1+s/s_{0})(1-\gamma s)^{2}]$ with $s_{0}=8.8\ \rm GeV^{2}$ were used for the case of proton Tomasi-Gustafsson et al. (2021); Tomasi-Gustafsson and Pacetti (2022), while the above dipole formula was used for neutron. In this work, we take the same formula for both proton and neutron. From the results shown in Figs. 3 (c) and (f), it is difficult to conclude the oscillation behaviour of the effective form factors of nucleon.  [^2] Same result is obtained if we used the three-pole formula for both proton and neutron. The dipole form was firstly obtained from the best fit to the electromagnetic form factors of proton in the space-like region. Nevertheless, the VMD model and the dipole form of $g(q^{2})$ can give a reasonable description of the experimental data on the baryon EMFFs at the considered energy region.
@@
Figure 3: The fitted results of total cross sections, effective form factors, and the oscillation part of the effective form factors for proton (left) and neutron (right). These error bands are obtained with the uncertainties of the fitted parameters. The red and green vertical dotted curves represent the $e^{+}e^{-}\to p\bar{p}$ and $n\bar{n}$ kinematic reaction threshold, respectively.
@@
Now we turn to the ratio of the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions. Until a few years ago, there had been some pioneering experimental efforts dedicated to the $e^{+}e^{-}\to n\bar{n}$ reaction by the FENICE experiment Antonelli et al. (1998). Even though the experimental data have large errors, and it is found that the total cross section of $e^{+}e^{-}\to n\bar{n}$ reaction is twice as large as the one of $e^{+}e^{-}\to p\bar{p}$ reaction, which is quite difficult to understand, since most theoretical predictions are opposite.
@@
Very recently, the BESIII Collaboration has published a new precise measurement of the total cross sections of $e^{+}e^{-}\to n\bar{n}$ reaction Ablikim et al. (2021b), and it is found that the ratio $R=\sigma_{e^{+}e^{-}\to p\bar{p}}/\sigma_{e^{+}e^{-}\to n\bar{n}}$ is larger than 1 at all the measured energies, which is contrast to the results obtained by the FENICE experiment. The new results of BESIII Collaboration indicates that the photon-proton coupling is stronger than the corresponding photon-neutron coupling. The SND Collaboration also measured the total cross sections of $e^{+}e^{-}\to n\bar{n}$ reaction with 2123 selected $n\bar{n}$ events for 14 energy points Achasov et al. (2022) and their experimental results are consistent with the results of BESIII Collaboration.
@@
Our theoretical results are shown in Fig. 4 by blue curve with error bands obtained from the uncertainties of the fitted parameters. The experimental data taken from BESIII and FENICE are also shown. Again, we can describe quite well the new experimental data by BESIII Collaboration Ablikim et al. (2021b).

### 29 Feb 2024 v2

@@ I Introduction
Nucleons (proton and neutron) are composite particles consisting of three valence quarks ($uud$ and $udd$) and a neutral sea of strong interaction. Despite the discovery of the proton over 100 years, an exact theoretical description of its internal structure has not been achieved within the framework of quantum chromodynamics (QCD) theory. This is due to the non-perturbative nature of QCD in the energy regime of the nucleon. On the theoretical side, the nucleon electromagnetic structure can be described by the electromagnetic form factors (EMFFs), which depends on squared of the four-momentum ($q^{2}$) of the exchanged virtual photon. The EMFFs in the space-like region ($q^{2}<0$) are real and they can be associated with the charge and magnetic distribution of the nucleons Pacetti et al. (2015). However, in the time-like region ($q^{2}>0$), the EMFFs become complex, providing information on the time evolution of the nucleon’s charge and magnetic moments at the nucleon-antinucleon pair formation point  Belushkin et al. (2007); Kuraev et al. (2012); Bianconi and Tomasi-Gustafsson (2015); Tomasi-Gustafsson et al. (2021). On the experimental side, the EMFFs of nucleons were measured by both the $e^{-}N\to e^{-}N$ elastic scattering Gayou et al. (2002, 2001); Andivahis et al. (1994); Bosted et al. (1992); Jones et al. (2000) and $e^{+}e^{-}\to N\bar{N}$ annihilation reactions Lees et al. (2013); Achasov et al. (2014); Ablikim et al. (2015); Akhmetshin et al. (2016); Ablikim et al. (2019a, 2020a, 2021a, 2021b); Huang and Ferroli (2021); Achasov et al. (2022). One can obtain the space-like form factors from the former process and the time-like form factors from the latter process, respectively.
@@
A comprehensive analysis of the nucleon’s electromagnetic form factors in both the space- and time-like regions using dispersion theory is conducted in Ref. Lin et al. (2022). For the periodic behaviour of the nucleon’s electromagnetic form factors, it has been studied in Refs. Cao et al. (2022); Xia et al. (2022); Dai et al. (2022); Bianconi and Tomasi-Gustafsson (2022); Yang et al. (2023); Milstein and Salnikov (2022); Tomasi-Gustafsson and Pacetti (2022); Pacetti et al. (2022); Qian et al. (2023); Dai et al. (2023), using various phenomenological methods. It is worthy mentioning that the periodic behaviour of the nucleon’s electromagnetic form factors are induced by those broad vector mesons according to the studies in Ref. Cao et al. (2022). However, these vector mesons couple to $N\bar{N}$ pair very weekly. In fact, a unified description of the time-like and space-like electromagnetic form factors of nucleons was proposed in Ref. Lomon and Pacetti (2012), where these low-lying vector mesons were taken into account. Meanwhile, in the time-like region, a revised Breit-Wigner formulas with momentum-dependent widths are needed for these vector resonances Lomon and Pacetti (2012). Indeed, the contributions of these vector mesons are important to the reactions of $e^{+}e^{-}$ annihilation into light hadrons Ablikim et al. (2019b, 2021c, 2021d, 2021e, c, 2020b, d, 2020c, 2021f).
@@
with $\gamma=1.41$ $\rm GeV^{-2}$ for the case of nucleon, and $s$ is the invariant mass square of the $e^{+}e^{-}$ annihilation process.
In this work, we study the $e^{+}e^{-}\to N\bar{N}$ reaction within the extended vector meson dominance model. Because proton and neutron are isospin doublets, the $p\bar{p}$ and $n\bar{n}$ states are expressed in terms of isospin 0 and 1 components. The mixtures of isovector and isoscalar for $p\bar{p}$ and $n\bar{n}$ of equal relative weight but different signs are imposed by the isospin symmetry as introduced by the underlying Clebsch-Gordan coefficients.In addition to these ground isovector $\rho$ meson and isoscalar $\omega$ meson, we will also study the important role played by the excited vector mesons with masses around 2.0 to 3.0 GeV,  [^1] It’s worth noting that due to the Okuba-Zweig-Iizuka rule Okubo (1963); Zweig (1964); Iizuka (1966), we do not consider the $\phi$ meson and its excited states here. with the aim of describing the new experimental data reported by BESIII Collaboration.
Since the information about the $\rho$ and $\omega$ excited states around $2.0$ to $3.0$ GeV is scarce Workman et al. (2022). Thus, it is necessary to rely on theoretical calculations. The BESIII Collaboration has also provided a large number of experimental data about $e^{+}e^{-}$ annihilation into light mesons Ablikim et al. (2019b, 2021c, 2021d, 2021e, c, 2020b, d, 2020c, 2021f), which provides the basis for our construction of light vector states. From the analyses of the $e^{+}e^{-}\to\pi^{+}\pi^{-}$, $\omega\pi^{0}$, and $\rho^{0}\eta^{\prime}$ reactions, these $\rho$ excited sates were studied in Refs. Wang et al. (2020); Zhou et al. (2022); Wang et al. (2021). The spectrum of excited $\rho$, $\omega$, and $\phi$ states were also investigated in Ref. Wang et al. (2022) using the modified Godfrey-Isgur model Song et al. (2015). By performing the phenomenological analysis, the two-body strong decay of the excited $\rho$ and $\rho_{3}$ states are systematically studied He et al. (2013). Based on the above theoretical calculations and the vector mesons listed in the particle data group (PDG) Workman et al. (2022), we take $\rho(2D)$, $\omega(3D)$, and $\omega(5S)$ into account in this work. Their masses and widths are collected in Table 1. By considering the contributions of these above low-lying vector meson excited states, we studied the $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions within the VMD model. It is found that the total cross sections of the two above reactions (corresponding to the nucleon effective form factors) and the nucleon electromagnetic form factors can be well described. However, the so-called oscillation behavior of the nucleon effective form factor discovered by the BESIII Collaboration Ablikim et al. (2021b) is not confirmed.
@@
It will be helpful to mention that the VMD model is a phenomenological approach and it works well in the low energy region. However, at high momentum transfer, the calculations from the VMD model are not consistent with that from the perturbative quantum chromodynamics Friedman (1991); Brodsky et al. (2017, 2018). In the present work, with the VMD model and the contributions from the low-lying excited $\omega$ and $\rho$ states, our calculations can give a reasonable description of the experimental measurements of the $e^{+}e^{-}\to N\bar{N}$ reaction in the considered energy region. Meanwhile, our calculation offers some important clues for the reaction mechanisms of the process $e^{+}e^{-}\to B\bar{B}$ and makes an effort to study the role of the vector states in relevant reactions.
This article is organized as follows: in the next section, we show the theoretical formalism of the nucleon EMFFs in the VMD model. Numerical results of the nucleon effective form factors, the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions, and the oscillating features of the nucleon EMFFs are shown in Sec. III. A short summary is given in the last section.
@@ II Theoretical formalism
with $\alpha=e^{2}/(4\pi)$ the fine-structure constant, $\tau=q^{2}/(4m_{N}^{2})$ and $\beta=\sqrt{1-4m_{N}^{2}/s}$ the phase-space factor. $m_{N}$ is the nucleon mass. Here, $C_{N}$ represents the S-wave Sommerfeld-Gamow factor accounts for the electromagnetic interaction of charged pointlike fermion pairs in the final state Arbuzov and Kopylova (2012). For proton, $C_{p}=y/(1-e^{-y})$ with $y=\frac{\alpha\pi}{\beta}\frac{2M_{p}}{\sqrt{s}}$, while for the neutron, $C_{n}=1$. Considering the $C_{N}$ factor, it is expected that the total cross section of $e^{+}e^{-}\to p\bar{p}$ reaction is nonzero at the reaction threshold.
@@
with $q^{2}=s$. Once we have $F_{1}(s)$ and $F_{2}(s)$, we can naturally obtain the electromagnetic form factors of the nucleon. Then the total cross sections of $e^{+}e^{-}\to N\bar{N}$ can be easily calculated.
Within the VMD model, the Dirac and Pauli form factors $F_{1}$ and $F_{2}$ for nucleon in the time-like region can be parameterized as,
@@
$B_{R}$ $=$ $\frac{m_{R}^{2}}{m_{R}^{2}-s-im_{R}\Gamma_{R}}.$ (10)
Here, $R\equiv\rho$, $\omega$, $\rho(2D)$, $\omega(3D)$, and $\omega(5S)$, and we take $m_{\rho}=775.26$ MeV, $\Gamma_{\rho}=147.4$ MeV, $m_{\omega}=782.66$ MeV, and $\Gamma_{\omega}=8.68$ MeV.
In addition, at $q^{2}=0$, with the constraints $G_{E}^{n}=0$ and $G_{M}^{n}=\mu_{n}$, $G_{E}^{p}=1$ and $G_{M}^{p}=\mu_{p}$, the coefficients $f_{1}^{n}$, $f_{2}^{n}$, $f_{1}^{p}$ and $f_{2}^{p}$ can be calculated,
@@
In the spacelike region, taking $Q^{2}=-q^{2}>0$, with the dipole form for $g(Q^{2})$, it is found that for the large value of $Q^{2}$, $F_{1}\sim\frac{1}{Q^{4}}$, and $F_{2}\sim\frac{1}{Q^{6}}$, which are consistent with the asymptotic behavior of $F_{1}$ and $F_{2}$ predicted by the perturbative quantum chromodynamics Brodsky and Farrar (1973); Lepage and Brodsky (1979, 1980).
@@ III NUMERICAL RESULTS
The results of the fitted parameters are listed in Table 2. The corresponding $\chi^{2}/\rm d.o.f$ is 1.6, where ${\rm d.o.f}$ is the number of dimension of the freedom. Note that the inclusion of more vector states will improve the fitting results, since it has more freedoms. We have explored such a possibility, but we have found some fitting problems. The fitted parameters have large uncertainties and strong correlations, which means that the inclusion of more vector states would not significantly improve the fitted results. Thus, we consider only the $\rho(2D)$, $\omega(3D)$, and $\omega(5S)$ states here.
@@
We show firstly the fitted numerical results of the electromagnetic form factors of proton and neutron in Fig. 1, comparing with the experimental data of the BESIII Collaboration. Within the contributions of the excited $\rho(2D)$, $\omega(3D)$, and $\omega(5S)$ states, we can obtain a good description of the electromagnetic form factors for both proton and neutron. For the electric form factor of proton, there is a dip structure around $2.3$ GeV which is because of the interference between the contributions of $\omega(3D)$ and $\omega(5S)$, and there is a clear peak for the $\omega(5S)$. This may indicate that, regarding to the current experimental data, the contribution of $\omega(5S)$ is necessary. It is expected that these theoretical results can be tested by future experiments.
Figure 1: Fitted results for the electromagnetic form factors of the proton and neutron comparing with the experimental data, which are taken from: BESIII 2020 Ablikim et al. (2020a) and BESIII 2023 Ablikim et al. (2023). The error bands are calculated with the uncertainties of the fitted parameters.
@@
Second, we show the fitted results of the modulus of the ratio $G_{E}/G_{M}$ for the proton in Fig. 2 compared with the experimental measurements. One can see that the experimental data have large errors. Again, there are dip and bump structures because of the contributions of of $\omega(3D)$ and $\omega(5S)$ states and their inteference. It is worthy mentioning that the ratio of $|G_{E}/G_{M}|$ equals to one at the reaction threshold as it should be. Clearly, more precise experimental data are needed to check our model calculations.
Figure 2: The ratio $|G_{E}/G_{M}|$ of proton comparing with the experimental data, which are taken from: $BABAR$ Lees et al. (2013), BESIII 2019 Ablikim et al. (2019a), BESIII 2021 Ablikim et al. (2021a), CMD3 Akhmetshin et al. (2016).
@@
The fitted numerical results of the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions are shown in Fig. 3 (a) and Fig. 4 (a), respectively. One can see that, thanks to the contribution of $\rho(2D)$, the platform behavior of the total cross section of $e^{+}e^{-}\to p\bar{p}$ can be well reproduced, and the experimental data of the $e^{+}e^{-}\to n\bar{n}$ can also be described within the fitted model parameters. Furthermore, these nonmonotonic structures around $2.2$ to $2.4$ GeV can be fairly well reproduced by considering the contributions of $\omega(3D)$ and $\omega(5S)$ states. In fact, a resonance with mass around 2300 MeV and width about 188 MeV is needed in the analysis of the total cross sections of $e^{+}e^{-}\to p\bar{p}$ reaction in Ref. Xia et al. (2022), where a simple Breit-Wigner amplitude for the $e^{+}e^{-}\to p\bar{p}$ reaction was taken.
@@
Since we can reproduce the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions, it is expected that the effective form factors of proton and neutron can also be described well, as shown in Fig. 3 (b) and Fig. 4 (b), respectively. This indicates that the nonmonotonic line shapes of the nucleon effective form factors can be explained within the VMD model, where the contributions of excited vector mesons are taken into account.
Before proceeding with further discussions, we need to clarify a few points. The excited vectors states of $\omega(3D)$ and $\omega(5S)$ can be replaced by the excited states of $\rho(3D)$ and $\rho(5S)$. This substitution does not significantly change the fitted results. As shown in Ref. Wang et al. (2021), the calculated masses of $\rho(3D)$ and $\rho(5S)$ are degenerate with $\omega(3D)$ and $\omega(5S)$, respectively, while the widths of $\rho(3D)$ and $\rho(5S)$ are $158$ and $80$ MeV, respectively.
Figure 3: The fitted results of total cross sections (a) and the effective form factors (b) for proton. The theoretical error bands are obtained with the uncertainties of the fitted parameters. The red vertical dotted curve represent the $e^{+}e^{-}\to p\bar{p}$ kinematic reaction threshold. The experimental data are taken from: $BABAR$ Lees et al. (2013), BESIII 2015 Ablikim et al. (2015), BESIII 2019 Ablikim et al. (2019a), BESIII 2020 Ablikim et al. (2020a), BESIII 2021 Ablikim et al. (2021a), and CMD3 Akhmetshin et al. (2016).
Figure 4: As in Fig. 3 but for the case of neutron. The experimental data are taken from: BESIII 2021 Ablikim et al. (2021b), SND 2014 Achasov et al. (2014), SND 2022 Achasov et al. (2022), and SND 2023 Achasov et al. (2023).
@@
Next, we will discuss the nonmonotonic features of the nucleon effective form factors. In general, the main part of the effective form factors of the nucleon can be described by a dipole form as shown in Eq. (1) with a global factor $c_{0}$, $G_{D}=c_{0}/(1-\gamma s)^{2}$. By fitting the experimental data of the effective form factors of proton and neutron, we get $c_{0}=5.54\pm 0.02$ for proton and $c_{0}=4.08\pm 0.04$ for neutron. Then, subtracting $G_{D}$ from the obtained effective form factor $|G_{\rm eff}|$, one can get the residual $G_{\rm osc}$,
@@
The predictions of our model for the oscillation parts of the proton and neutron effective form factors are shown in Fig. 5 (a) and (b), respectively. It is found that the model predictions are in agreement with the subtracted data obtained with Eq. (15). One the other hand, for the oscillation part, we have performed new fitting with two formulas: $G_{\mathrm{osc}}=A\cos(C\sqrt{s}+D)/(1-\gamma s)^{2}$ Dai et al. (2022) and $G_{\rm{osc}}=A\exp(-Bp)\cos(Cp+D)$ Ablikim et al. (2021b) with $p$ the relative momentum between $N$ and $\bar{N}$ in the final state. The first one is formally referred to as Model I, while the other is designated as Model II. The fitted results are also shown in Fig. 5 (a) and (b). One can see that, with the dipole form of $G_{\rm{D}}$, the oscillation part of proton can not be well described, and the two fitting results are almost the same. While for the neutron, the oscillation part can be fairly well reproduced, and the two fittings are very similar.
@@
Note that, in Ref. Ablikim et al. (2021b), to get the conclusion which the effective form factor of both proton and neutron show a periodic behavior, different formula of $G_{D}$ were taken for proton and neutron. A three-pole formula $F_{3p}=F_{0}/[(1+s/s_{0})(1-\gamma s)^{2}]$ with $s_{0}=8.8\ \rm GeV^{2}$ were used for the case of proton Tomasi-Gustafsson et al. (2021); Tomasi-Gustafsson and Pacetti (2022), while the above dipole formula was used for neutron. In this work, we take the same formula for both proton and neutron. From the results shown in Fig. 5, it is difficult to conclude the oscillation behaviour of the effective form factors of nucleon.  [^2] Same result is obtained if we used the three-pole formula for both proton and neutron. The dipole form was firstly obtained from the best fit to the electromagnetic form factors of proton in the space-like region. Nevertheless, the VMD model and the dipole form of $g(q^{2})$ can give a reasonable description of the experimental data on the baryon EMFFs at the considered energy region.
@@
Figure 5: The fitted results of the oscillation part of the effective form factors for proton (a) and neutron (b). These error bands of ”This work” are obtained with the uncertainties of the fitted parameters. The red and green vertical dotted curves represent the $e^{+}e^{-}\to p\bar{p}$ and $n\bar{n}$ kinematic reaction threshold, respectively. The experimental data are taken from: $BABAR$ Lees et al. (2013), BESIII 2015 Ablikim et al. (2015), BESIII 2019 Ablikim et al. (2019a), BESIII 2020 Ablikim et al. (2020a), CMD3 Akhmetshin et al. (2016), BESIII 2021 Ablikim et al. (2021a, b), SND 2014 Achasov et al. (2014), SND 2022 Achasov et al. (2022), and SND 2023 Achasov et al. (2023).
@@
Now we turn to the ratio of the total cross sections of $e^{+}e^{-}\to p\bar{p}$ and $e^{+}e^{-}\to n\bar{n}$ reactions. Until a few years ago, there had been some pioneering experimental efforts dedicated to the $e^{+}e^{-}\to n\bar{n}$ reaction by the FENICE experiment Antonelli et al. (1998). Even though the experimental data have large errors, and it is found that the total cross section of $e^{+}e^{-}\to n\bar{n}$ reaction is twice as large as the one of $e^{+}e^{-}\to p\bar{p}$ reaction, which is quite difficult to be understood, since most theoretical predictions are opposite. For example, the ratio of $R=\sigma_{e^{+}e^{-}\to p\bar{p}}/\sigma_{e^{+}e^{-}\to n\bar{n}}$ is $7/3$ predicted by the SU(6) symmetric nucleon wave function Farrar and Jackson (1975) and it is the naive $e^{2}_{u}/e^{2}_{d}=4$ from the quark charge ratio in the constituent quark model Chernyak and Zhitnitsky (1984).
@@
Very recently, the BESIII Collaboration has published a new precise measurement of the total cross sections of $e^{+}e^{-}\to n\bar{n}$ reaction Ablikim et al. (2021b), and it is found that the ratio $R=\sigma_{e^{+}e^{-}\to p\bar{p}}/\sigma_{e^{+}e^{-}\to n\bar{n}}$ is larger than 1 at all the measured energies, which is contrast to the results obtained by the FENICE experiment. The new results of BESIII Collaboration indicates that the photon-proton coupling is stronger than the corresponding photon-neutron coupling. The SND Collaboration also measured the total cross sections of $e^{+}e^{-}\to n\bar{n}$ reaction for 14 energy points Achasov et al. (2022) and their experimental results are consistent with the results of BESIII Collaboration.
@@
Our theoretical results of the ratio $R$ are shown in Fig. 6 by blue curve with error bands obtained from the uncertainties of the fitted parameters. The experimental data taken from BESIII and FENICE are also shown. Again, we can describe quite well the new experimental data by BESIII Collaboration Ablikim et al. (2021b).